# On the generalization of language models from in-context learning and finetuning: a controlled study

### v1

@@ Abstract
Large language models exhibit exciting capabilities, yet can show surprisingly narrow generalization from finetuning — from failing to generalize to simple reversals of relations they are trained on, to missing logical deductions that can be made from trained information. These failures to generalize from fine-tuning can hinder practical application of these models. However, language models’ in-context learning shows different inductive biases, and can generalize better in some of these cases. Here, we explore these differences in generalization between in-context- and fine-tuning-based learning. To do so, we constructed several novel datasets to evaluate and improve models’ ability to generalize from finetuning data. The datasets are constructed to isolate the knowledge in the dataset from that in pretraining, to create clean tests of generalization. We expose pretrained large models to controlled subsets of the information in these datasets — either in context, or through fine-tuning — and evaluate their performance on test sets that require various types of generalization. We find overall that in data-matched settings, in-context learning can generalize more flexibly than fine-tuning (though we also find some qualifications of prior findings, such as cases when fine-tuning can generalize to reversals embedded in a larger structure of knowledge). We build on these findings to propose a method to enable improved generalization from fine-tuning: adding in-context inferences to finetuning data. We show that this method improves generalization across various splits of our datasets and other benchmarks. Our results have implications for understanding the inductive biases of different modes of learning in language models, and practically improving their performance.
@@ Introduction
Language models (LMs) pretrained on huge corpora of internet text become efficient in-context learners; they can generalize from a few examples of a task to answer new instances (Brown et al., 2020; Team et al., 2023). Pretrained LMs can also be fine-tuned for downstream tasks using relatively few examples—though achieving good generalization from fine-tuning often requires hundreds to thousands of examples (e.g. Kirstain et al., 2022; Vieira et al., 2024). Indeed, the generalization from fine-tuning on a particular example can be surprisingly limited; for example, LMs fine-tuned on a statement like “B’s mother is A” fail to generalize to answer “Who is A’s son?” (Berglund et al., 2024). However, LMs can readily answer questions about these types of reverse relations in context (e.g. Lampinen et al., 2024b). Do in-context learning and fine-tuning therefore result in different patterns of generalization (cf. Chan et al., 2022b; Shen et al., 2023; Russin et al., )? If so, how does this affect how we should adapt models to downstream tasks? In this paper, we explore these questions.
To do so, we construct controlled synthetic datasets of factual knowledge. We design these datasets to have complex and self-consistent structures, but avoid any overlap with knowledge that may be present in pretraining corpora. We create train and test splits of these datasets that involve different types of generalization: reversals or chaining multiple logical inferences into a syllogism. We then evaluate how well large pretrained language models generalize to these test sets through fine-tuning, or through in-context learning—via placing the entire training set (or large subsets thereof) in context. We also explore various methods of improving generalization, such as data augmentation. Overall, we find that across various datasets, in-context learning (ICL) generalizes better than finetuning, however, finetuning generalization can be improved, and in fact made better than ICL (on the original data), by spending more train time compute to augment the training dataset using in-context inferences.
@@
  • We also propose a fine-tuning method that breaks up correlations among sentences, thereby improving the benefits of dataset augmentation.
@@
We evaluate learning and generalization across several different datasets which are crafted to respectively isolate features of the generalization problem, or situate it within a broader set of learning challenges. We also draw on datasets from prior work.
@@
Simple reversals: Each training example consists of a brief context (e.g. “Did you know that”) followed by a single sentence containing a comparison of two entities, e.g. “femp are more dangerous than glon.” There are a hundred such facts provided (with the dimension of comparison for each sampled from a set of 28 features, e.g. “brighter”, “heavier,” etc.), each of which is repeated across 10 different training articles with a randomly sampled context statement in each. The test set consists of forced choices between the correct reversal and a contradictory relation, e.g.: “glon are less/more dangerous than femp.”
Simple syllogisms: There are 69 training examples, each of which contains a piece of context (which entities are referred to) and two statements that form a syllogism. For example: “The relationships between glon, troff and yomp are: All glon are yomp. All troff are glon.” The test examples test whether the model makes a correct inference from that syllogistic form, by providing a context of two entities from the conclusion, and then scoring all possible statements about their relationship involving the quantifiers and relations in the dataset. Following Lampinen et al. (2024b), we omit the partial negative form (some X are not Y) pattern; thus, there are six such possible statements (the product of “all”, “some”, and “no” together with the two directions each of those relations could take). For example, the context for the test example corresponding to the syllogism above would be “The relationships between yomp and troff are:” and the correct answer would be “All troff are yomp.”
@@
We use the reversal dataset proposed by Berglund et al. (2024) that contains one-sentence descriptions of fictional celebrities. Dataset examples can have the celebrity name (e.g. ‘Daphne Barrington’) preceding the description (e.g. ‘the director of "A Journey Through Time."’) or vice-versa. During finetuning, the training dataset is presented in one order, say where the name precedes the description, and then tested on the reversed order, where the name follows the description.
@@
The use of nonsense terms does, however, potentially pose tokenization challenges; however, we ameliorate these by generating short 4-5 letter nonsense words using plausible combinations of phonemes for English (sampled via Gao et al., 2023). Moreover, we show in the next sections that in practice the models can readily make inferences about these nonsensical entities in context, so use of unfamiliar words is not the primary bottleneck on model performance.
@@
  • Rephrasings of trained facts that don’t change the direction: used for validation that the material is learned.
  • Reversals of trained facts.
@@
  • Category holdouts: only one fact about a category is present in training: what its parent category is. All possible inferences from that fact are tested. This overlaps with some aspects of the syllogism splits, except that the information about the target category is strictly more limited, thus limiting other cues that could aid generalization.
@@ Methods
Our fine-tuning experiments mainly involve tuning Gemini 1.5 Flash (Team et al., 2024) on our datasets. We generally fine-tune with batch size 8 or 16 and learning rate $3\cdot 10^{-4}$ for 200-1000 steps (depending on dataset size and loss).
@@
To perform full-dataset in-context evaluation, we concatenate the documents in a training dataset and provide them as context to the (instruction tuned) model. We then prompt the model to use that context to answer the question. We randomly subsample the documents by a factor of 8x when evaluating models in-context on the largest datasets, as there is some redundancy in the datasets, and we find that models experience interference as the context length scales.
@@
Our key approach to dataset augmentation is to take advantage of the in-context generalization of models in order to improve the coverage of the fine-tuning dataset. We approach this via several methods, but all chiefly aim at the goal of spending training time compute for in-context inference to create more finetuning data, in order to improve generalization out-of-context at test time.
@@
Some datasets, such as the fictional celebrities dataset by Berglund et al. (2024) and our semantic structure dataset, contain documents comprising multiple sentences. We find that splitting these documents at sentence-level, into multiple finetuning examples, leads to a much improved finetuning performance – even after accounting for the total dataset size and gradient steps. We explore two ways of splitting a document into sentences 1) Independent Splitting: wherein all sentences are split independently into separate training examples, 2) Cumulative Splitting: wherein a document is split such that for any target sentence in an example, all preceding sentences of the document are added to the context of that example. For an $n$-sentence document, both methods produce $n$ training examples, although with different context sizes. We analyze the effect of sentence splitting on the model generalization in appendix B.1. In the following sections, we assume independent sentence splitting unless stated otherwise.
@@
In Fig. 2, we first explore the generalization issues in the context of the reversal curse phenomenon and dataset released by Berglund et al. (2024). We replicate the authors’ finding that finetuning on the forward directions does not produce generalization to the reversals. The authors of that work noted briefly that in-context the models could perform better at this task. We study this more systematically by presenting the entire dataset in context, and find that the model performs nearly at ceiling on the reversals—thus providing a strong demonstration of the benefits of in-context learning over finetuning. Finetuning with data augmented with in-context inferences yields similarly high test performance. Simple finetuning, on the other hand, has near zero accuracy as the finetuned model always prefers those (incorrect) celebrity name completions that were seen as targets during training, regardless of the context. Finally, a pretrained model performs near chance on the test set, indicating a lack of contamination.
@@
We then correspondingly test the models on the simple nonsense version of our reversal dataset (cf. Fig 3). We find a weaker, but still noticeable, benefit of ICL over finetuning in this setting, and again stronger benefits of augmented finetuning. The difference in benefits compared to the experiments above may be due to differences in the plausibility of the relationships in question, e.g. the possibility that nonsense words interfere to some degree with the model’s reasoning over longer contexts (see Appx. B.2).
@@
We next test the simple syllogisms dataset (cf. Fig 3). Again, the pretrained model performs at chance, indicating a lack of contamination. Finetuning on the dataset does yield some non-trivial above-chance generalization; perhaps because for certain flavors of syllogism the logical inferences are compatible with simpler linguistic patterns — e.g., tuning on sequences like “All X include Y.” “All Y include Z.” might make the model more likely to predict “All X include Z” from pure associations. However, following simpler associative patterns is not valid for most syllogistic forms, e.g. if the universal quantifier is replaced by the existential in the examples above. Perhaps because of this, in-context learning yields much stronger performance; using in-context learning to augment the training dataset further improves performance overall.
@@
We then test the broader semantic structure benchmark (Fig. 4) which integrates multiple kinds of generalization into a richer learning setting. In this setting we evaluate performance on: a) rephrased training statements (that preserve the direction of the relation) denoted as ‘train’ in the figure, b) reversals, c) syllogistic inferences, and d) propositions about held out categories. Across these settings, we find an overall benefit of ICL over finetuning, though the magnitude of this benefit varies depending on the split. We find some improvements in generalization to even rephrased information from training, and more dramatic improvements on reversals and syllogisms. However, the category-holdout split remains difficult, and improvements from ICL are minimal. Furthermore, we continue to find that augmenting the finetuning data with in-context inferences improves performance from finetuning, in many cases outperforming even ICL. (N.B., in this setting we do not use sentence splitting for the finetuning baseline, as we find it impairs performance and we want to compare to the strongest baselines; results with sentence-splitting for the baseline, along with other ablations, can be found in Appx. B.4.)
@@
Figure 4: On the more richly structured semantic dataset, in-context learning still moderately outperforms finetuning. Furthermore, augmentation continues to show some benefit — even in rephrased questions about trained facts that do not involve reversal. However some generalization splits, such as the category-level holdout, remain very challenging. (Error bars are standard errors computed over task subsets featuring different types of the inferences in question, e.g. reversals of property relations vs. reversals of category inclusion relations.)
@@
We also performed some preliminary experiments exploring learning and generalization of process-type (rather than semantic) knowledge. Specifically, we focused on executing a simple pseudo-mathematical procedure which is a function to function transform similar to taking a derivative, but with uncommon rules. We created a train-test compositional generalization split in which certain combinations of rules are seen in training, but other combinations are held out for testing. The overall pattern of results is generally consistent with those above—with ICL outperforming finetuning, and augmentations improving finetuning performance. However, process knowledge requires distinct evaluation methods, and the effects correspondingly appear to be driven by distinct factors compared to the semantic knowledge experiments above; thus, we present the experiments and results in full detail in Appendix B.5.
@@
In-context learning: A variety of works have explored in-context learning (Brown et al., 2020; Dong et al., 2022; Lampinen et al., 2024a). Several works have studied patterns in the learning and generalization of ICL, either empirically (Garg et al., 2022; Chan et al., 2022a), mechanistically (Olsson et al., 2022), or theoretically (Xie et al., 2023; Zhang et al., 2024). Several recent works have found that even scaling to hundreds or thousands of in-context examples can improve LM performance on challenging datasets (Agarwal et al., 2024; Anil et al., 2024; Bertsch et al., 2024). Our work similarly relies on the ability of models to learn from many documents in context in order to generate the inferences for augmentation.
@@
Data augmentation: A wide variety of works have explored how LLMs can be used to augment data for improved performance from small or narrow datasets, e.g. to improve cross-lingual generalization (e.g. Whitehouse et al., 2023). Ding et al. (2024) reviews some of this broader literature. There have also been targeted attempts to fix specific issues like the reversal curse with hardcoded augmentations (Golovneva et al., 2024). A closely related work by Akyürek et al. (2024) proposes “deductive closure training” to improve coverage via prompting language models to generate deductive inferences from training documents. Padmanabhan et al. (2024) similarly propose generating continuations and then distilling them into the model. Several concurrent works Chen et al. (2024); Ruan et al. (2025) suggest that having LMs generate additional reasoning directions, and using these to augment their training data, can improve performance on reasoning tasks. Our experiments show that in controlled settings, without the possibility of dataset contamination, similar approaches to augmenting small finetuning datasets can yield improved generalization—and relate the performance to that achieved through more basic finetuning and in-context learning.
Synthetic data: An equally broad literature has explored applications of synthetic data to improving LM performance; see Liu et al. (2024) for a recent survey. Earlier works considered hand-designed synthetic data targeted to improve generalization using domain knowledge of areas like linguistics or mathematics (e.g. Pratapa et al., 2018; Wu et al., 2021). More recent approaches have focused on generating the data directly from language models, either filtering with ground-truth scores (Zelikman et al., 2022) or simply through self-consistency (Wang et al., 2023; Huang et al., 2023). While a recent prominent article argued that training on synthetic data leads models to collapse (Shumailov et al., 2024), other works have noted that in settings where synthetic data is mixed with the original data performance continues to improve (Gerstgrasser et al., 2024). We correspondingly find that in our setting, rather than being harmful, synthesizing augmented data with the models consistently improves performance (even on rephrased information from the training split in the semantic structure dataset). These results contribute to the ongoing discussion of how incorporating synthetic data affects model performance.

### v2

@@ Abstract
Large language models exhibit exciting capabilities, yet can show surprisingly narrow generalization from finetuning. E.g. they can fail to generalize to simple reversals of relations they are trained on, or fail to make simple logical deductions based on trained information. These failures to generalize from fine-tuning can hinder practical application of these models. On the other hand, language models’ in-context learning shows different inductive biases, and can generalize better in some cases. Here, we explore these differences in generalization between in-context- and fine-tuning-based learning. To do so, we constructed several novel datasets to evaluate and improve models’ abilities to generalize from finetuning data. The datasets are designed to create clean tests of generalization, by isolating the knowledge in the dataset from that in pretraining. We expose pretrained large models to controlled subsets of the information in these datasets — either in context, or through fine-tuning — and evaluate their performance on test sets that require various types of generalization. We find overall that in data-matched settings, in-context learning can generalize more flexibly than fine-tuning (though we also find some qualifications of prior findings, such as cases when fine-tuning can generalize to reversals embedded in a larger structure of knowledge). We build on these findings to propose a method to enable improved generalization from fine-tuning: adding in-context inferences to finetuning data. We show that this method improves generalization across various splits of our datasets and other benchmarks. Our results have implications for understanding the inductive biases of different modes of learning in language models, and practically improving their performance.
@@ Introduction
Language models (LMs) pretrained on huge corpora of internet text become efficient in-context learners; they can generalize from a few examples of a task to answer new instances (Brown et al., 2020; Gemini Team Google, 2023). Pretrained LMs can also be fine-tuned for downstream tasks using relatively few examples—though achieving good generalization from fine-tuning often requires hundreds to thousands of examples (e.g. Kirstain et al., 2022; Vieira et al., 2024). Indeed, the generalization from fine-tuning on a particular example can be surprisingly limited; for example, LMs fine-tuned on a statement like “B’s mother is A” fail to generalize to answer “Who is A’s son?” (Berglund et al., 2024; cf. Allen-Zhu and Li, 2025). However, LMs can readily answer questions about these types of reverse relations in context (e.g. Lampinen et al., 2024b). Furthermore, transformers show different inductive biases when generalizing from weights vs from context (cf. Chan et al., 2022b; Shen et al., 2023; Russin et al., ). How do patterns of generalization differe between in-context learning and fine-tuning? What are the implications for how we should adapt models for new tasks or information? In this paper, we explore these questions.
To do so, we construct controlled synthetic datasets of factual knowledge. We design these datasets to have complex and self-consistent structures, while avoiding any overlap with knowledge that may be present in pretraining corpora. We create train and test splits of these datasets that involve different types of generalization: reversals or chaining multiple logical inferences into a syllogism. We then evaluate how well large pretrained language models generalize to these test sets through fine-tuning, or through in-context learning—via placing the entire training set (or large subsets thereof) in context. We also explore various methods of improving generalization, such as data augmentation. Overall, we find that across various datasets, in-context learning (ICL) generalizes better than finetuning. However, finetuning generalization can be improved, and in fact can even exceed ICL generalization, by spending more train time compute to augment the training dataset using in-context inferences.
@@
  • We also propose a fine-tuning method that breaks up correlations among sentences, amplifying the benefits of augmentation.
@@
We evaluate learning and generalization across several datasets that are crafted to isolate different aspects of generalization, or to situate it within a broader set of learning challenges. We also draw on datasets from prior work.
@@
Simple reversals: Each training example consists of a brief preamble (e.g. “Did you know that”) followed by a single statement comparing two entities, e.g. “femp are more dangerous than glon.” There are a hundred such comparisons provided (with the comparison words sampled from a set of 28, e.g. “brighter”, “heavier,”). Each comparison is repeated across 10 different training examples, each time paired with a randomly sampled preamble. The test set consists of forced choices between the correct reversal and a contradictory relation, e.g.: “glon are less/more dangerous than femp.”
Simple syllogisms: There are 69 training examples. Each example consists of a preamble (which specifies the entities under discussion), followed by two statements that form a syllogism. For example: “The relationships between glon, troff and yomp are: All glon are yomp. All troff are glon.” The test examples evaluate whether the model makes a correct inference from that syllogistic form. They provide a preamble specifying two entities from the training example, and then score all possible statements about the relationship between the two entities involving the quantifiers and relations in the dataset. Following Lampinen et al. (2024b), we omit the partial negative form (some X are not Y); thus, there are six such possible statements (the product of “all”, “some”, and “no” together with the two directions each of those relations could take). For example, the preamble for the test example corresponding to the syllogism above would be “The relationships between yomp and troff are:” and the correct answer would be “All troff are yomp.”
@@
We use the reversal dataset proposed by Berglund et al. (2024) that contains one-sentence descriptions of fictional celebrities. Dataset examples can have the celebrity name (e.g. ‘Daphne Barrington’) preceding the description (e.g. ‘the director of "A Journey Through Time."’) or vice-versa. During finetuning, the training dataset is presented in one order, e.g. where the name precedes the description, and then tested on the reversed order, where the name follows the description.
@@
The use of nonsense terms potentially poses tokenization challenges; however, we ameliorate these by generating short 4-5 letter nonsense words using plausible combinations of phonemes for English (sampled via Gao et al., 2023). Moreover, we show in the next sections that in practice the models can readily make inferences about these nonsensical entities in context, so use of unfamiliar words is not the primary bottleneck on model performance.
@@
  • Rephrasings of trained facts that don’t change the direction of wording: used for validation that the material is learned.
  • Reversals of trained facts, i.e. the two entities are in reverse order, and the relational words are adjusted accordingly.
@@
  • Category holdouts: only one fact about a category is present in training: what its parent category is. All possible inferences from that fact are tested. This overlaps with some aspects of the syllogism splits, except that the information about the target category is more strictly limited, thus limiting other cues that could aid generalization.
@@ Methods
Our fine-tuning experiments mainly involve tuning Gemini 1.5 Flash (Gemini Team Google, 2024) on our datasets. We generally fine-tune with batch size 8 or 16 and learning rate $3\cdot 10^{-4}$ for 200-1000 steps (depending on dataset size and loss).
@@
To perform full-dataset in-context evaluation, we concatenate the documents in a training dataset and provide them as context to the (instruction-tuned) model. We then prompt the model to use that context to answer the question. We randomly subsample the documents by a factor of 8x when evaluating models in-context on the largest datasets, as there is some redundancy in the datasets, and we find that models experience interference as the context length scales.
@@
Our key approach to dataset augmentation is to take advantage of the in-context generalization of models in order to improve the coverage of the fine-tuning dataset. We approach this via several methods, but all chiefly aim at the goal of spending training time compute for in-context inference to create more finetuning data, in order to improve generalization out-of-context (i.e., without requiring additional information in context) at test time.
@@
Some datasets, such as the fictional celebrities dataset by Berglund et al. (2024) and our semantic structure dataset, contain documents comprising multiple sentences that are logically or semantically linked. We find that splitting these documents at sentence-level, into multiple finetuning examples, leads to a much improved finetuning performance – even after accounting for the total dataset size and gradient steps. We explore two ways of splitting a document: 1) Independent Splitting: wherein all $n$ sentences of a document are split independently into $n$ separate training examples, 2) Cumulative Splitting: wherein an $n$-sentence document is split into $n$ examples, with the $i$th example contains all sentences up to the $i$th sentence. We analyze the effect of sentence splitting on the model generalization in appendix B.1. In the following sections, we assume independent sentence splitting unless stated otherwise.
@@
In Fig. 2, we first explore generalization in the context of the reversal curse phenomenon and dataset released by Berglund et al. (2024). We replicate the finding that finetuning on the forward directions does not produce generalization to the reversals. The authors of that work noted briefly that in-context the models could perform better at this task. We study this more systematically by presenting the entire dataset in context, and find that the model performs nearly at ceiling on the reversals—thus providing a strong demonstration of the benefits of in-context learning over finetuning. Finetuning with data augmented with in-context inferences yields similarly high test performance. Simple finetuning, on the other hand, has near zero accuracy as the finetuned model always prefers those (incorrect) celebrity name completions that were seen as targets during training, regardless of the context. Finally, a pretrained model performs near chance on the test set, indicating a lack of contamination.
@@
We then correspondingly test the models on the simple nonsense version of our reversal dataset (Fig 3, left). We find a weaker, but still noticeable, benefit of ICL over finetuning in this setting, and again stronger benefits of augmented finetuning. The difference in benefits compared to the experiments above may be due to differences in the plausibility of the relationships in question, e.g. the possibility that nonsense words interfere to some degree with the model’s reasoning over longer contexts (see Appx. B.2).
@@
We next test the simple syllogisms dataset (Fig 3, right). Again, the pretrained model performs at chance, indicating a lack of contamination. Finetuning on the dataset does yield some non-trivial above-chance generalization; perhaps because for certain flavors of syllogism the logical inferences are compatible with simpler linguistic patterns — e.g., tuning on sequences like “All X include Y.” “All Y include Z.” might make the model more likely to predict “All X include Z” from pure associations. However, following simpler associative patterns is not valid for most syllogistic forms, e.g. if the universal quantifier is replaced by the existential in the examples above. Perhaps because of this, in-context learning yields much stronger performance; using in-context learning to augment the training dataset further improves performance overall.
@@
We then test the semantic structure benchmark (Fig. 4), which integrates multiple kinds of generalization in a richer learning setting. In this setting, we evaluate performance on: a) rephrased training statements (that preserve the direction of the relation) denoted as ‘train’ in the figure, b) reversals, c) syllogistic inferences, and d) propositions about held out categories. Across these settings, we find an overall benefit of ICL over finetuning, though the magnitude of this benefit varies depending on the split. We find some improvements in generalization to even rephrased information from training, and more dramatic improvements on reversals and syllogisms. However, the category-holdout split remains difficult, and improvements from ICL are minimal. Furthermore, we continue to find that augmenting the finetuning data with in-context inferences improves performance from finetuning, in many cases outperforming even ICL. (N.B., in this setting we do not use sentence splitting for the finetuning baseline, as we find it impairs performance and we want to compare to the strongest baselines; results with sentence-splitting for the baseline, along with other ablations, can be found in Appx. B.4.)
@@
Figure 4: On the more richly structured semantic dataset, in-context learning still moderately outperforms finetuning. Furthermore, augmentation continues to show benefit for generalization from finetuning — even in rephrased questions about trained facts that do not involve reversal (top left). However some generalization splits, such as the category-level holdout, remain very challenging. (Error bars are standard errors computed over task subsets featuring different types of the inferences in question, e.g. reversals of property relations vs. reversals of category inclusion relations.)
@@
We also performed some preliminary experiments exploring learning and generalization of process-type (rather than semantic) knowledge, i.e. the ability to apply novel procedures to inputs. Specifically, we focused on executing a simple pseudo-mathematical procedure that we call “derivatoids”, which transforms math-like expressions in a manner similar to computing a derivative. We created a train-test compositional generalization split in which certain combinations of rules are seen in training (again, either in context or via finetuning), but other combinations are held out for testing. The overall pattern of results is generally consistent with those above. ICL outperforms finetuning, especially in the low-shot regime, and augmentations improve finetuning performance. However, process knowledge requires distinct evaluation methods, and the effects correspondingly appear to be driven by distinct factors compared to the semantic knowledge experiments above; thus, we present the experiments and results in full detail in Appendix B.5.
@@
In-context learning: A variety of works have explored in-context learning (Brown et al., 2020; Dong et al., 2022; Lampinen et al., 2024a). Several works have studied patterns in the learning and generalization of ICL, either empirically (Garg et al., 2022; Chan et al., 2022a), mechanistically (Olsson et al., 2022), or theoretically (Xie et al., 2023; Zhang et al., 2024; Elmoznino et al., 2024). Several recent works have found that even scaling to hundreds or thousands of in-context examples can improve LM performance on challenging datasets (Agarwal et al., 2024; Anil et al., 2024; Bertsch et al., 2024). Our work similarly relies on the ability of models to learn from many documents in context in order to generate the inferences for augmentation.
@@
Factual learning & generalization: A variety of works have studied how language models learn and generalize factual information, both in pretraining (e.g. Allen-Zhu and Li, 2024, 2025; Zucchet et al., 2025) and during finetuning (e.g. Berglund et al., 2024; Gekhman et al., 2024). Many of these works have focused on failures to generalize like the “Reversal Curse” (Berglund et al., 2024), while others looked into increases in hallucination as LLMs are pretrained (Zucchet et al., 2025) or finetuned on new knowledge (Gekhman et al., 2024). However, these works have generally not performed data-matched experiments comparing the generalization of in-context learning and finetuning.
Data augmentation: A wide variety of works have explored how LLMs can be used to augment data for improved performance from small or narrow datasets, e.g. to improve zero-shot task performance (Maini et al., 2024) or to improve cross-lingual generalization (e.g. Whitehouse et al., 2023). Ding et al. (2024) reviews some of this broader literature. There have also been targeted attempts to fix specific issues like the reversal curse with hardcoded augmentations (Golovneva et al., 2024). A closely related work by Akyürek et al. (2024) proposes “deductive closure training” to improve coverage via prompting language models to generate deductive inferences from training documents. Padmanabhan et al. (2024) similarly propose generating continuations and then distilling them into the model. Several concurrent works Chen et al. (2024); Ruan et al. (2025) suggest that having LMs generate additional reasoning directions, and using these to augment their training data, can improve performance on reasoning tasks. Yang et al. (2024) use language models to extract entities in the training documents, and generate synthetic data reasoning about links between these entities; like our results, they find this method of “rearranging” knowledge in synthetic data helps to improve downstream performance. Our experiments show that in controlled settings, without the possibility of dataset contamination, similar approaches to augmenting small finetuning datasets can yield improved generalization—and relate the performance to that achieved through more basic finetuning and in-context learning.
Synthetic data: An equally broad literature has explored other applications of synthetic data to improving LM performance; see Liu et al. (2024) for a recent survey. Earlier works considered hand-designed synthetic data targeted to improve generalization using domain knowledge of areas like linguistics or mathematics (e.g. Pratapa et al., 2018; Wu et al., 2021). More recent approaches have focused on generating the data directly from language models, either filtering with ground-truth scores (Zelikman et al., 2022) or simply through self-consistency (Wang et al., 2023; Huang et al., 2023). While a recent prominent article argued that training on synthetic data leads models to collapse (Shumailov et al., 2024), other works have noted that in settings where synthetic data is mixed with the original data performance continues to improve (Gerstgrasser et al., 2024). We correspondingly find that in our setting, rather than being harmful, synthesizing augmented data with the models consistently improves performance (even on rephrased information from the training split in the semantic structure dataset). These results contribute to the ongoing discussion of how incorporating synthetic data affects model performance.